OpenStudy (anonymous):

Find the slope of line passing through the points (5, 9) and (5, -2) A. -11 B. 0 C. 11 D. undefined

jimthompson5910 (jim_thompson5910):

Change in y: 9 - (-2) = 9 + 2 = 11 Change in x: 5 - 5 = 0 Slope = (change in y)/(change in x) = 11/0 = undefined So the slope is undefined

jimthompson5910 (jim_thompson5910):

this line is completely vertical

jimthompson5910 (jim_thompson5910):

All vertical lines have undefined slopes

OpenStudy (anonymous):

so d????